Budget Total
(in Billions of Dollars)
2006 2007 2008 2009 2010 2011 2012 2013 2014
Revenues 2,407 2,568 2,524 2,105 2,162 2,304 2,450 2,775 3,032
Outlays 2,655 2,729 2,983 3,518 3,456 3,603 3,537 3,455 3,523
Surplus (deficit) (248) (161) (459) (1,413) (1,294) (1,301) (1,087) (680) (492)


From 2008 to 2010, by how much did each of the following change?

Instructions: Enter your responses rounded to the nearest whole number. Indicate negative changes with a (-) negative sign.

(a) Tax revenue: $  billion

(b) Government spending (outlays): $  billion

(c) The budget deficit: $  billion
Answer #1

Solution:-

(A). Tax Revenue

The tax revenue of 2008 to 2010 can be calculated as follows:-

Tax Revenue = Tax Revenue 2010 - Tax Revenue 2008

= 2,162 - 2,524

= -362

The tax revenue is decrease by 362 Billion

(B). Government Spending

The government spending of 2008 to 2010 can be calculated as follows:-

Government spending = Outlays 2010 - Outlays 2008

   = 3,456 - 2,983

= 473

Government spending is increase by 473 Billion.

(C). Budget Deficit:-

The budget deficit of 2008 to 2010 can be calculated as follows:-

Budget Deficit = Deficit 2010 - Deficit 2008

= -1294 - (-459)

= -835

The budget deficit decrease by 835 billion.